if a recipe calls for 7 cups of rice and you use a half of teaspoon of salt per cup, how much salt would you use?
Answer by josmiceli(19441)

( 1/2 tspn/cup ) x ( 7 cups ) = 7/2 tspns of salt
7/2 = 3 1/2 tspns